OpenStudy (anonymous):

Help? Lighthouse A, located on an ocean shore, uses a 10-kW light. Lighthouse B uses a 20-kW light and is located 8 km out to sea from a point 6 km down the beach from Lighthouse A. 2. A man is walking down the beach away from lighthouse A and toward point C. When he is x kilometers away from lighthouse A and has not yet reached point C, write the illumination he receives as a function of x in simplest form.

OpenStudy (anonymous):

okay im not uderstanding where point c is. is it the intersection at the ocean shore?

OpenStudy (anonymous):

One sec.

OpenStudy (anonymous):

okay give me a minute

OpenStudy (anonymous):

Alright!

OpenStudy (anonymous):

im sorry i dont think i can get this one if someone else answers it i would like to know the answer

OpenStudy (anonymous):

haha No problem. :) Thanks for trying!

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@dumbcow @phi

OpenStudy (anonymous):

@ganeshie8

OpenStudy (dumbcow):

you can measure the distance from point B using pythagorean thm \[d = \sqrt{8^{2} + (6-x)^{2}} = \sqrt{x^{2}-12x+100}\] is illumination measured by watts*distance ? if so then function is \[l = 10x+20\sqrt{x^{2}-12x+100}\]